OpenStudy (anonymous):

There are 30 students on the school's student council. A special homecoming dance committee is to be formed by randomly selecting 6 students from student council. How many possible committees can be formed?

Parth (parthkohli):

30C6, because we're making 6 combinations out of 30.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {30! \over (30 - 6)!6!}}\) Calculator needed alert

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {30 \times 29 \times 28 \times 26 \times25 \over 6 \times 5 \times 4 \times 3 \times 2 \times 1}}\)